What is the present value of a maturity value of $2,484 bearing interest at 14% for 90days? Use formula P=M/1 + rt
1
Expert's answer
2011-05-12T11:28:30-0400
M = P(1+rt) =
& = 2484 *(1+ 0.14 * 90/360) =
& = 2570.94